# Gradient-based Sample Selection for Faster Bayesian Optimization

### v1

@@ 4.2 Gradient-based Sample Selection
As the number of observed samples increases, fitting a GP model can become prohibitively expensive, especially in large-scale scenarios. A common remedy is to work with a subset of samples of size $M\ll N$, thereby reducing the computational cost of GP updates. Once the kernel parameters are fixed, the efficiency and effectiveness of GP model fitting in BO are closely related to the quality of this chosen subset. This raises the question: How do we choose a subset that remains representative and informative?
Inspired by the success of gradient-based subset selection methods in machine learning, we propose leveraging gradient information to guide the selection of such subsets within BO. To this end, we introduce a gradient-based sample selection methodology to ensure representativeness within a limited sample buffer size. By harnessing gradient information, our approach maintains a carefully chosen subset of samples that not only eases computational burdens, but also preserves model quality, even as the sample set size grows.
We begin by modeling the objective function $f$ with a GP and setting a buffer size $M$. Initially, the algorithm observes $f$ at $n_{0}$ samples, retaining these initial samples to preserve global information critical to the model. After each subsequent evaluation, if the number of samples exceeds $M$, we perform a gradient-based sample selection step to ensure that only $M$ representative samples are kept for the next GP update.
@@ 4.3 Gradient-based Sample Selection BO
Detailed Implementations. In the initialization phase, $n_{0}$ initial samples $\{(x_{i},y_{i})\}_{i=1}^{n_{0}}$ are observed; and the initial sample set $D$, the buffer size $M$, and total budget $N$ are specified. In each iteration, the GP posterior is updated based on the current sample set $D$; an acquisition function (e.g., UCB) is built to select the next point $x_{t}$, and the corresponding observation $y_{t}=f(x_{t})$ is obtained; and the sample $(x_{t},y_{t})$ is added to $D$. To manage each iteration’s computational complexity, a buffer check and gradient-based selection step are performed. Specifically, if the current size of $D$ is less than or equal to $M$, the GP is updated using all samples in $D$. Otherwise, a gradient-based sample selection step is performed to identify a set of the most representative samples. Note that the $n_{0}$ initial samples and the newly acquired sample $(x_{t},y_{t})$ are always added into the subset, as they provide base information for the GP model and ensure that recently observed information is always retained, respectively. Besides the initial $n_{0}$ and the newly observed samples, $(M-n_{0}-1)$ samples are selected by minimizing the sum of pairwise cosine similarities among their gradients. The resulting subset $\mathcal{U}$, containing $M$ samples, is then used to fit the GP model. The complete procedure is outlined in Algorithm 4.3.
@@
(1) Dynamic Buffer Size. In practice, the buffer size should be prespecified by the users. However, the value is often unavailable in advance. Instead, we propose a dynamic adjustment mechanism to determine the buffer size. We define a tolerable maximum factor $Z$ to accelerate GP computations. Let $\bar{T}$ be the average evaluation time for a single sample point estimated based on the initial iterations and $T_{\text{current}}$ be the current iteration’s computation time. If $T_{\text{current}}$ exceeds the user-specified threshold $Z\times\bar{T}$, the buffer size is set to be the number of all current samples, i.e., $M=|D|$. This adaptive strategy ensures that the algorithm balances computational efficiency with the goal of utilizing as much data as possible, thereby maintaining high predictive accuracy without incurring excessive costs.
@@
(2) Preserving Initial Samples and Latest Observations. During the procedure, the initial $n_{0}$ samples are always included in the subset. These samples capture essential information about the overall function landscape, and good initialization samples give us useful information about the global situation. Additionally, the newly acquired sample, $(x_{t},y_{t})$, is also included in the subset. This ensures that the GP model incorporates the latest data, maintaining its relevance and accuracy. Consequently, the algorithm prevents valuable information from being prematurely excluded. Additionally, this essentially alleviates a limitation of sparse GP in BO (McIntire et al., 2016): the constrained representation size may hinder the full integration of new observations into the model.
@@
(3) Random Perturbation to Escape Local Optima. Subset selection may cause the algorithm to become trapped in the local optima, as it iteratively selects a subset of “locally optimal” samples. This issue arises when certain samples are frequently chosen by the acquisition function but consistently excluded from the sample subset, hindering the exploration of other valuable samples. To address this issue, Gaussian noise $\epsilon_{i}\sim\mathcal{N}(0,\sigma^{2})$ is added to each computed gradient $g_{i}$, resulting in perturbed gradients $\tilde{g}_{i}=g_{i}+\epsilon_{i}$. This perturbation introduces variability into the gradient-based selection process, helping the algorithm escape local optima and facilitating the exploration of a more diverse set of samples. The parameter $\sigma$ serves as a tunable control for the degree of perturbation.
@@
1:  Initialization: Place a GP prior on $f$. Obtain $n_{0}$ initial samples $D=\{(x_{i},y_{i})\}_{i=1}^{n_{0}}$. Set the buffer size $M>n_{0}$, total budget $N$, and average initial iteration time $\bar{T}$.
@@
3:     Update posterior $p(f\mid D)$.
4:     Select $x_{t}=\arg\max_{x}\alpha(x;p(f|D))$, where $\alpha$ is the acquisition function.
5:     Evaluate $y_{t}=f(x_{t})$ and set $D=D\cup\{(x_{t},y_{t})\}$.
6:     if $|D|>M$ then
7:        Compute gradients $g_{i}$ for samples $(x_{i},y_{i})\in D$.
8:        Add the newest sample $(x_{t},y_{t})$ into the subset.
9:        Form a subset $\mathcal{U}$ by forcing in the $n_{0}$ initial samples and the newest sample $(x_{t},y_{t})$.
10:        Select $(M-n_{0}-1)$ samples from $D$ such that the sum of $\{\tilde{g}_{i}\}$ is minimized.
11:        Update the GP using the $M$ selected samples.
12:     else
13:        Update the GP using all samples in $D$.
14:     end if
15:     Let $T_{\text{current}}$ be the current iteration time.
16:     if $T_{\text{current}}>Z\times\bar{T}$ then
17:        Set $M=|D|$.
@@ 5 Theoretical Analysis
Gaussian Process Upper Confidence Bound (GP-UCB (Srinivas et al., 2009)) is a popular algorithm for sequential decision-making problems. We propose an extension to GP-UCB by incorporating gradient-based sampling. In this section, we analyze the error of the subset fitted GP and prove that the regret of the GSSBO with GP-UCB algorithm is bounded.
@@
(Error in the Subset-Fitted GP) This theorem establishes bounds on the difference between the posterior mean and variance under a subset fitted GP approximation and those of the full set fitted GP. Given a GP with kernel matrix $\mathbf{K}_{\mathcal{D}\mathcal{D}}$, a low-rank approximation $\hat{\mathbf{K}}=\mathbf{K}_{\mathcal{D}\mathcal{U}}\mathbf{K}_{\mathcal{U}\mathcal{U}}^{-1}\mathbf{K}_{\mathcal{U}\mathcal{D}}$ constructed from $M$ inducing samples and a test sample $\mathbf{x}_{*}$, the posterior predictive mean and variance errors satisfy:
$|\Delta\mu(\mathbf{x}_{*})|\leq\|\mathbf{k}_{*\mathcal{D}}\|\|\mathbf{y}\|C_{1}\|\mathbf{K}_{\mathcal{D}\mathcal{D}}-\hat{\mathbf{K}}\|,$
$|\Delta\sigma^{2}(\mathbf{x}_{*})|\leq\|\mathbf{k}_{*\mathcal{D}}\|^{2}C_{1}\|\mathbf{K}_{\mathcal{D}\mathcal{D}}-\hat{\mathbf{K}}\|,$
where $\mathbf{k}_{*\mathcal{D}}\in\mathbb{R}^{N}$ means the covariance vector between the test sample $\mathbf{x}_{*}$ and all training samples in $\mathcal{D}$, $C_{1}=\|({\mathbf{K}_{\mathcal{D}\mathcal{D}}+\sigma_{n}^{2}\mathbf{I}})^{-1}\|\|({\hat{\mathbf{K}}+\sigma_{n}^{2}\mathbf{I}})^{-1}\|$.
@@
Assumption 5.1. Assume there exist constants $a$, $b$, and $L$ such that the kernel function $k(\mathbf{x},\mathbf{x}^{\prime})$ satisfies a Lipschitz continuity condition, providing confidence bounds on the derivatives of the GP sample paths $f$:
$P\left(\sup_{\mathbf{x}\in\mathcal{X}}\left|\frac{\partial f}{\partial x_{j}}\right|>L\right)\leq ae^{-L^{2}/b^{2}}\quad\text{for }j=1,\ldots,p.$
A typical example of such a kernel is the squared exponential kernel $k(\mathbf{x},\mathbf{x}^{\prime})=\sigma^{2}\exp\left(-\frac{\|\mathbf{x}-\mathbf{x}^{\prime}\|^{2}}{2l^{2}}\right)$, where $l$ is the length-scale parameter and $\sigma^{2}$ represents the noise variance. Then we propagate the error in Theorem 5.1 through the GP posterior to bound the GSSBO regret.
@@
(Regret Bound for GSSBO with UCB) Let $\mathcal{X}\subseteq[0,r]^{p}$ be compact and convex, $p\in\mathbb{N},r>0$, let $A=\|\mathbf{k}_{*\mathcal{D}}\|\|\mathbf{y}\|\;C_{1}\,(\lambda_{M+1}+\epsilon\sum_{i=1}^{N}(\mathbf{K}_{\mathcal{D}\mathcal{D}})_{ii}^{2})$, $\delta\geq A/\sigma_{\min}$. Under Assumption 1, for any arbitrarily small $\delta\in(0,1)$, choose
$\beta_{n}=2\log\frac{4\pi_{n}}{\delta}+{2p}\log\left(n^{2}brp\sqrt{\log\left(\frac{4pa}{\delta}\right)}\right)-\frac{A}{\sigma_{\min}},$
where $\sum_{n\geq 1}\pi_{n}^{-1}=1,\pi_{n}>0$. As $n\to\infty$, we obtain a regret bound of $\mathcal{O}^{*}\left(\sqrt{pN\gamma_{N}}\right)$. Specifically, with $C=\frac{8}{\log(1+\sigma^{-2})},$ we have:
$P\left(R_{N}\leq\sqrt{CN\beta_{N}\gamma_{N}}\right)\geq 1-\delta.$
(Two-Phase Regret) Let $N$ be the total number of rounds. In the first $M_{\text{initial}}$ rounds, one applies the full GP-UCB. Subsequently, from round $M+1$ to $N$, one switches to the gradient-based subset strategy. The total regret satisfies $R_{N}\;=\;R_{M}^{(\text{full})}\;+\;R_{N-M}^{(\text{selected})},$ where $R_{M}^{(\text{full})}\leq\sqrt{\,C\,M\,\beta_{M}\,\gamma_{M}+2}$ and $R_{N-M}^{(\text{selected})}\leq\sqrt{\,C\,(N-M)\,\beta_{(N-M)}\,\gamma_{(N-M)}}$.
The sketch proof for the main theorem is relegated to the Appendix.
The main theoretical challenge lies in evaluating the error between the low-rank approximation $\hat{\mathbf{K}}$ and the full $\mathbf{K}_{\mathcal{D}\mathcal{D}}$. By invoking spectral norm inequalities and the Nyström approximation theory, $\|\mathbf{K}_{\mathcal{D}\mathcal{D}}-\hat{\mathbf{K}}\|$ can be bounded. Then, we merge the resulting linear and $\beta$-scaled error terms into a single penalty in the UCB construction. This establishes a regret bound for GSSBO, which is similar to that of classical GP-UCB. From a practical relevance perspective, Theorem 5.1 indicates that limiting the GP to a smaller, well-chosen subset does not substantially degrade posterior accuracy in either the mean or variance estimates. Restricting the subset size $M$ confers significant computational savings while ensuring performance closely matches that of a standard GP-UCB using all samples. Of note, we also observe that, compared with the classical UCB results, our GSSBO retains the same fundamental structure of an upper confidence bound approach. Still, it restricts the GP fitting to a gradient-based sample subset, lowering computational costs.
@@ 6 Experiments
Figure 2: Cumulative regret of algorithms on the Eggholder2, Hart6, Levy20, Powell50, Rastrigin100 functions and NAS experiment.
In this section, we conduct numerical experiments to illustrate the superior efficiency of GSSBO. The objective of the numerical experiments are threefold: (1) to evaluate computational efficiency; (2) to assess optimization performance; and (3) to validate the theoretical analysis.
@@
We choose UCB as the acquisition function in GSSBO, and compare GSSBO with the following benchmarks: (1) Standard GP-UCB (Srinivas et al., 2009), which retains all observed samples without any selection; (2) Random Sample Selection GP-UCB (RSSBO), which mirrors our approach in restricting the sample set size but chooses which samples to keep purely at random; (3) VecchiaBO (Jimenez & Katzfuss, 2023), which utilizes the Vecchia approximation method to condition the GP likelihood on the nearest neighbors in a predefined maximization order; (4) SVIGP (Hensman et al., 2013), which applies stochastic variational inference by optimizing pseudo-points to approximate the GP posterior; and (5) LR-First m (Williams & Seeger, 2000), which uses a low-rank approximation based on the first $m$ samples in a maximization order, with the conditioning set fixed across all samples. We employ a Matérn $5/2$ kernel for the GP, with hyperparameters learned via maximum likelihood estimation. Both GSSBO and RSSBO use the same buffer size $M$, dynamically adjusted by a parameter $Z=4$. The gradient-perturbation noise is set to $\sigma^{2}=0.01$. Each experiment is repeated 50 times to rule out accidental variations, and the total budget is 400.

### v2

@@ 4.2 Gradient-based Sample Selection
As the number of observed samples increases, fitting a GP model can become prohibitively expensive, especially in large-scale scenarios. A common remedy is to work with a subset of samples of size $M\ll N$, thereby reducing the computational cost of GP updates. The efficiency and effectiveness of GP model fitting in BO are closely related to the quality of the chosen subset. This raises the question: How do we choose a subset that remains representative and informative? Inspired by the success of gradient-based subset selection methods in machine learning, we propose leveraging gradient information to guide the selection of such subsets within BO. To this end, we introduce a gradient-based sample selection methodology to ensure representativeness within a limited sample buffer size. By harnessing gradient information, our approach maintains a carefully chosen subset of samples that not only eases computational burdens, but also preserves model quality, even as the sample set size grows. We begin by modeling the objective function $f$ with a GP and setting a buffer size $M$. Initially, the algorithm observes $f$ at $n_{0}$ samples, retaining these initial samples to preserve global information critical to the model. After each subsequent evaluation, if the number of samples exceeds $M$, we perform a gradient-based sample selection step to ensure that only $M$ representative samples are kept for the next GP update.
@@ 4.3 Gradient-based Sample Selection BO
Detailed Implementations. In the initialization phase, $n_{0}$ initial samples $\{(x_{i},y_{i})\}_{i=1}^{n_{0}}$ are observed; and the initial sample set $D$, the buffer size $M$, and total budget $N$ are specified. In each iteration, the GP posterior is updated on the current sample set $D$, an acquisition function (e.g., UCB) is built to select the next point $x_{t}$, the corresponding observation $y_{t}=f(x_{t})$ is obtained, and the sample $(x_{t},y_{t})$ is added to $D$. To manage each iteration’s computational complexity, a buffer check and gradient-based selection step are performed. Specifically, if the current size of $D$ is less than or equal to $M$, the GP is updated using all samples. Otherwise, a gradient-based sample selection step is performed to identify a set of the most representative samples. Note that the $n_{0}$ initial samples and the newly acquired sample $(x_{t},y_{t})$ are always added into the subset, as they provide base information for the GP model and ensure that recently observed information is always retained, respectively. Besides the initial $n_{0}$ and the newly observed samples, $(M-n_{0}-1)$ samples are selected by minimizing the sum of pairwise cosine similarities among their gradients. The resulting subset $\mathcal{U}$, containing $M$ samples, is then used to fit the GP model. The complete procedure is outlined in Algorithm 4.3.
@@
(1) Dynamic Buffer Size. In practice, the buffer size should be prespecified by the users. However, the value is often unavailable in advance. Instead, we propose a dynamic adjustment mechanism to determine the buffer size. We define a tolerable maximum factor $Z$ to accelerate GP computations. Let $\bar{T}$ be the average time for a single initial iteration and $T_{\text{current}}$ be the current iteration’s computation time. If $T_{\text{current}}$ exceeds the user-specified threshold $Z\times\bar{T}$, the buffer size is set to be the number of all current samples, i.e., $M=|D|$. This adaptive strategy ensures that the algorithm balances computational efficiency with the goal of utilizing as much data as possible, thereby maintaining high predictive accuracy without incurring excessive costs.
@@
(2) Preserving Initial Samples and Latest Observations. During the procedure, the initial $n_{0}$ samples are always included in the subset. These samples capture essential information about the overall function landscape, and good initialization samples give us useful information about the global situation. Additionally, the newly acquired sample, $(x_{t},y_{t})$, is also included in the subset. This ensures that the GP model incorporates the latest data, maintaining its relevance and accuracy. Consequently, the algorithm prevents valuable information from being prematurely excluded. Additionally, this essentially alleviates a limitation of sparse GP in BO [17]: the constrained representation size may hinder the full integration of new observations into the model.
@@
(3) Random Perturbation to Escape Local Optima. Subset selection may cause being trapped in the local optima, as it iteratively selects a subset of “locally optimal” samples. This issue arises when certain samples are frequently chosen by the acquisition function but consistently excluded from the sample subset, hindering the exploration of other valuable samples. To address this issue, Gaussian noise $\epsilon_{i}\sim\mathcal{N}(0,(\sigma_{d})^{2})$ is added to each computed gradient $g_{i}$, resulting in perturbed gradients $\tilde{g}_{i}=g_{i}+\epsilon_{i}$. This perturbation introduces variability into the selection process, helping the algorithm escape local optima and facilitating the exploration of a more diverse set of samples.
Here we highlight the difference between our method and SparseGP. SparseGP methods require an additional correction vector or term to compensate for the diagonal discrepancies between the low-rank kernel matrix and the high-rank kernel matrix. In contrast, our method leverages gradient information to select the most representative samples, thereby constructing a low-rank approximation that directly approximates the full kernel matrix without the need for extra correction vectors. This not only simplifies the model structure but also reduces the additional computational overhead. The overall speed of GSSBO is often faster than Sparse GP, which requires multiple rounds of optimization for variational inference.
@@
1:  Initialization: Obtain $n_{0}$ initial samples $D=\{(x_{i},y_{i})\}_{i=1}^{n_{0}}$, and fit an initial GP model. Set buffer size $M>n_{0}$, total budget $N$, average initial iteration time $\bar{T}$, and threshold factor $Z$. Initialize switched $\leftarrow$ false.
@@
3:     Select $x_{t}=\arg\max_{x}\;\alpha\bigl{(}x;\,p(f\mid D)\bigr{)},$ where $\alpha$ is the acquisition function.
4:     Evaluate $y_{t}=f(x_{t})$ and set $D\leftarrow D\cup\{(x_{t},y_{t})\}$.
5:     if not switched then
6:        Let $T_{\text{current}}$ be the current iteration time.
7:        if $T_{\text{current}}>Z\times\bar{T}$ then
[…]
9:        end if
10:     end if
11:     if switched then
12:        Compute gradients $g_{i}$ for all $(x_{i},y_{i})\in D$.
13:        Form subset $\mathcal{U}$ by forcing in the $n_{0}$ initial samples and the newest sample $(x_{t},y_{t})$.
14:        Select the remaining $M-n_{0}-1$ samples from $D$ via the gradient-based criterion.
15:        Update the GP using the $M$ samples in $\mathcal{U}$.
16:     else
17:        Update the GP using all samples in $D$.
@@ 5 Theoretical Analysis
Gaussian Process Upper Confidence Bound (GP-UCB [46]) is a popular algorithm for sequential decision-making problems. We propose an extension to GP-UCB by incorporating gradient-based sampling. Some assumptions and notations follow the previous proof. Here we analyze the error of the subset fitted GP and prove that the regret of the GSSBO with GP-UCB algorithm is bounded.
@@
(Error in the Subset-Fitted GP) This theorem establishes bounds on the difference between the posterior mean and variance under a subset fitted GP approximation and those of the full set fitted GP. For a noisy sample $\mathbf{y}=f(\mathbf{X})+\bm{\epsilon},~{}\bm{\epsilon}\sim\mathcal{N}(0,\sigma_{n}^{2}\mathbf{I}),$i.i.d. Given a GP with kernel matrix $\mathbf{K}_{\mathcal{D}\mathcal{D}}$, a low-rank approximation $\hat{\mathbf{K}}=\mathbf{K}_{\mathcal{D}\mathcal{U}}\mathbf{K}_{\mathcal{U}\mathcal{U}}^{-1}\mathbf{K}_{\mathcal{U}\mathcal{D}}$ constructed from $M$ inducing samples and a test sample $\mathbf{x}_{*}$, the posterior predictive mean and variance errors satisfy:
$|\Delta\mu(\mathbf{x}_{*})|\leq\|\mathbf{k}_{*\mathcal{D}}\|\|\mathbf{y}\|C_{M}\|\mathbf{K}_{\mathcal{D}\mathcal{D}}-\hat{\mathbf{K}}\|,|\Delta\sigma^{2}(\mathbf{x}_{*})|\leq\|\mathbf{k}_{*\mathcal{D}}\|^{2}C_{M}\|\mathbf{K}_{\mathcal{D}\mathcal{D}}-\hat{\mathbf{K}}\|,$ (4)
where $\mathbf{k}_{*\mathcal{D}}\in\mathbb{R}^{N}$ means the covariance vector between the test sample $\mathbf{x}_{*}$ and all training samples in $\mathcal{D}$, $C_{M}=\|({\mathbf{K}_{\mathcal{D}\mathcal{D}}+\sigma_{n}^{2}\mathbf{I}})^{-1}\|\|({\hat{\mathbf{K}}+\sigma_{n}^{2}\mathbf{I}})^{-1}\|$.
@@
Assumption 5.1. Assume there exist constants $a$, $b$, and $L$ such that the kernel function $k(\mathbf{x},\mathbf{x}^{\prime})$ satisfies a Lipschitz continuity condition, providing confidence bounds on the derivatives of the GP sample paths $f$: $P\left(\sup_{\mathbf{x}\in\mathcal{X}}\left|\frac{\partial f}{\partial x_{j}}\right|>L\right)\leq ae^{-L^{2}/b^{2}}\quad\text{for }j=1,\ldots,d.$ A typical example of such a kernel is the squared exponential kernel $k(\mathbf{x},\mathbf{x}^{\prime})=\sigma^{2}\exp\left(-\frac{\|\mathbf{x}-\mathbf{x}^{\prime}\|^{2}}{2l^{2}}\right)$, where $l$ is the length-scale parameter and $\sigma^{2}$ represents the noise variance. Then we propagate the error in Theorem 5.1 through the GP posterior to bound the GSSBO regret.
@@
[…]
$P\left(R_{N}\leq\sqrt{C_{1}{T_{N}}\beta_{{T_{N}}}\gamma_{T_{N}}}\right)\geq 1-\delta.$ (5)
(Two-Phase Regret) Let ${T_{N}}$ be the total number of rounds. In the first ${T_{M}}$ rounds, one applies the full GP-UCB. Subsequently, from round ${T_{M+1}}$ to ${T_{N}}$, one switches to the gradient-based subset strategy. The total regret satisfies $R_{T_{N}}\;=\;R_{{T_{M}}}^{(\text{full})}\;+\;R_{T_{N-M}}^{(\text{selected})},$ where $R_{{T_{M}}}^{(\text{full})}\leq\sqrt{\,C_{1}\,{T_{M}}\,\beta_{T_{M}}\,\gamma_{T_{M}}}+2$ and $R_{T_{N-M}}^{(\text{selected})}\leq\sqrt{\,C_{1}\,(T_{N}-T_{M})\,\beta_{T_{N-M}}\,\gamma_{T_{N-M}}}$.
The sketch proof for the main theorem is relegated to the Appendix. The main theoretical challenge lies in evaluating the error between the low-rank approximation $\hat{\mathbf{K}}$ and the full $\mathbf{K}_{\mathcal{D}\mathcal{D}}$. By invoking spectral norm inequalities and the Nyström approximation theory, $\|\mathbf{K}_{\mathcal{D}\mathcal{D}}-\hat{\mathbf{K}}\|$ can be bounded. We prove that select $\mathcal{U}$ by maximizing gradient-direction diversity coincides with the greedy Nyström column-selection, then derive select $\mathcal{U}$ by maximizing gradient-direction diversity is better than random selection in standard Nyström approximation. Then, we merge the resulting linear and $\beta$-scaled error terms into a single penalty in the UCB construction. We also determine the smallest subset size $M_{\min}$. This establishes a regret bound for GSSBO, which is similar to that of classical GP-UCB. From a practical relevance perspective, Theorem 5.1 indicates that limiting the GP to a smaller, well-chosen subset does not substantially degrade posterior accuracy in either the mean or variance estimates. Restricting the subset size $M$ confers significant computational savings while ensuring performance closely matches that of a standard GP-UCB using all samples. Of note, we also observe that, compared with the classical UCB results, our GSSBO retains the same fundamental structure of an upper confidence bound approach. Still, it restricts the GP fitting to a gradient-based sample subset, lowering computational costs.
@@ 6 Experiments
In this section, we conduct numerical experiments to illustrate the superior efficiency of GSSBO. The objective of the numerical experiments is threefold: (1) to evaluate computational efficiency; (2) to assess optimization performance; and (3) to validate the theoretical analysis. To assess the performance of our proposed methods, we test five benchmark functions, Eggholder2, Hart6, Levy20, Powell50, Rastrigin100 and Neural Architecture Search (NAS).
@@
We choose UCB as the acquisition function in GSSBO, and compare GSSBO with the following benchmarks: (1) Standard GP-UCB [46]; (2) Random Sample Selection GP-UCB (RSSBO), which mirrors our approach in restricting the sample set size but chooses which samples randomly; (3) VecchiaBO [12], which utilizes the Vecchia approximation method in BO; (4) LR-First m [14], which uses a low-rank approximation based on the first $m$ samples. (5) SparseGP [15], which introduces a small number of inducing points to obtain a sparse approximation of GP; (6) IPA [24], which multiplies an expected-improvement-based quality function by the GP posterior and selects inducing points regarding quality-diversity. and (7) mini-META [25] scaling GP optimization by repeatedly evaluating each selected point until its posterior uncertainty falls below a preset threshold. (More baseline can be found in the appendix.) We employ a Matérn $5/2$ kernel for the GP, with hyperparameters learned via maximum likelihood estimation. Both GSSBO and RSSBO use the same buffer size $M$, dynamically adjusted by a parameter $Z=4$, $M$ is usually around 100 and will change with $Z$. The gradient-perturbation noise is set to $\sigma^{2}=0.01$. The size of the initial set is 20. Each experiment is repeated 50 times, and the total number of iterations is 1000. All experiments were conducted on a MacBook Pro with Apple M2 Pro (10-core CPU, 16 GB unified RAM).